# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 14.14 Type ${\bf G}_{2}$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,6$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G_{2}/{B})$. In other words, the above collection is obtained from the collection Lemma 14.37 by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection above is right dual to that of Lemma 14.37.
Proof.
According to Proposition (7.5), one needs to check the following orthogonalities:
  1. [1)]
  2. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-6],\mathcal{L}(-\rho))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-6],{{\sf A}_{i}})=0$ for $i\neq-6$
  3. 2.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{{\sf A}_{i}})=0$ for $i\neq-5$
  4. 3.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{{\sf A}_{i}})=0$ for $i\neq-5$
  5. 4.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{{\sf A}_{i}})=0$ for $i\neq-4$
  6. 5.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{{\sf A}_{i}})=0$ for $i\neq-4$
  7. 6.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  8. 7.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  9. 8.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\beta})[-2],{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\beta})[-2],\Psi_{2}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\omega_{\beta})[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  10. 9.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})[-2],{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})[-2],\Psi_{2}^{\omega_{\alpha}}))={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  11. 10.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\alpha}})=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  12. 11.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{-\alpha}))\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{\beta}})=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{-\alpha}))\otimes\mathcal{L}(\rho)[-1],,\Psi_{1}^{\omega_{\alpha}})={\sf k}$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(\omega_{-\alpha}))\otimes\mathcal{L}(\rho)[-1],,{{\sf A}_{i}})=0$ for $i\neq-1$
  13. 12.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{L}(0))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),{{\sf A}_{i}})=0$ for $i\neq 0$.
[…]
2): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\alpha}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})))={\sf k}$. By Remark 14.39, (6), the group in question is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(X,(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\beta}-\rho)[5])$; tensoring the sequence (57) for $\omega_{\alpha}$ with $\mathcal{L}(-\omega_{\beta}-\rho)$, the coboundary homomorphism of the associated long exact cohomology sequence gives the required isomorphism as $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-2\rho))={\sf k}[-6]$.
3): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{\beta}}\otimes\mathcal{L}(\rho)[-5],{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$. By Remark 14.39, (4), the group in question is isomorphic to $\mathop{\mathrm{R\Gamma}}\nolimits(X,(\Psi_{1}^{\omega_{\beta}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}-\rho)[5])$ and we obtain the required isomorphism as in the previous item 2).
4): $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))={\sf k}$. By Remark 14.39, (3), the group in question is isomorphic to $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{2}^{\omega_{\alpha}},\mathcal{U}_{2}\otimes\mathcal{L}(-\rho)[4])$. Applying $\mathop{\mathrm{Hom}}\nolimits_{X}(-,\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))$ to (89) and using isomorphisms $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))=0$, $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\omega_{\alpha}),\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))=0$, and $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(2\omega_{\alpha}),\mathcal{U}_{2}\otimes\mathcal{L}(-\rho))=\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-2\rho))=\sf k[-6]$ (use (91) for the last isomorphism), we obtain the required isomorphism.
5): $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(\rho)[-4],(\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}))={\sf k}$. By Serre duality, the group in question is isomorphic to $\mathop{\mathrm{RHom}}\nolimits_{X}((\Psi_{1}^{\omega_{\alpha}})^{\ast}\otimes\mathcal{L}(-\omega_{\alpha}),{\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(-\rho)[2])^{\ast}=\mathop{\mathrm{RHom}}\nolimits_{X}((\Psi_{1}^{\omega_{\alpha}})^{\ast},{\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(-\omega_{\beta})[2])^{\ast}$. Now using the sequences (57) and Proposition 14.22, one calculates the group $\mathop{\mathrm{RHom}}\nolimits_{X}((\Psi_{1}^{\omega_{\alpha}})^{\ast},{\sf D}_{\alpha}\Psi_{1}^{\omega_{\alpha}}\otimes\mathcal{L}(-\omega_{\beta}))$ to be isomorphic to $\sf k[-2]$; this gives the required isomorphism.
6): $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(\rho)[-3],{\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}}))={\sf k}$. By Remark 14.39, ${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=\sf S$ and ${\sf D}_{\alpha}(\Psi_{2}^{\omega_{\alpha}})=\mathcal{E}$. To compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{E})$, use the sequence (95) from which one obtains $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{E})=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\Lambda^{2}\mathcal{U}_{2}^{\perp})$ as $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{L}(0)=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}^{\ast}\otimes\mathcal{L}(-\rho))=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf D}_{\beta}{\sf S}^{\ast}\otimes\mathcal{L}(-\rho))=\mathop{\mathrm{R\Gamma}}\nolimits(Y_{\beta},{\pi_{\beta}}_{\ast}{\sf S}^{\ast}\otimes{\pi_{\beta}}_{\ast}\mathcal{L}(-\rho))=0$. Now $\Lambda^{2}\mathcal{U}_{2}^{\perp}$ has a resolution which is the second exterior power of the tautological short exact sequence $0\rightarrow\mathcal{U}_{2}^{\ast}\rightarrow\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(0)\rightarrow\mathcal{U}_{2}^{\perp}\rightarrow 0$:
$0\rightarrow{\sf S}^{2}\mathcal{U}_{2}^{\ast}\rightarrow\mathcal{U}_{2}^{\ast}\otimes\nabla_{\omega_{\alpha}}\rightarrow\Lambda^{2}\nabla_{\omega_{\alpha}}\otimes\mathcal{L}(0)\rightarrow\Lambda^{2}\mathcal{U}_{2}^{\perp}\rightarrow 0$ (96)
Apply $\mathop{\mathrm{Hom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],-)$ to the above sequence. We have already obtained $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{L}(0))=0$. A straightforward calculation with the help of the dual to sequence (91) and Theorem 14.1 shows that $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{U}_{2}^{\ast})=0$. We have to compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],{\sf S}^{2}\mathcal{U}_{2}^{\ast})=\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}^{2}\mathcal{U}_{2}^{\ast}\otimes{\sf S}^{\ast}\otimes\mathcal{L}(-\rho)[3])$. Now the weights of the $G$-equivariant bundle ${\sf S}^{2}\mathcal{U}_{2}^{\ast}$ are equal to $2\omega_{\beta}-2\omega_{\alpha},\omega_{\beta},2\omega_{\alpha}$ and using also the isomorphism ${\sf S}={\sf S}^{\ast}\otimes\mathcal{L}(-\omega_{\alpha})$, we obtain three cohomology groups (ignoring the shift by $[3]$):
$\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}\otimes\mathcal{L}(\omega_{\beta}-2\omega_{\alpha})),\ \mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}),\ \mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}\otimes\mathcal{L}(2\omega_{\alpha}-\omega_{\beta}))$
Now both groups $\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S})$ and $\mathop{\mathrm{R\Gamma}}\nolimits(X,{\sf S}\otimes\mathcal{L}(2\omega_{\alpha}-\omega_{\beta}))$ are trivial. Calculating the weights of the remaining $G$-equivariant bundle ${\sf S}\otimes\mathcal{L}(\omega_{\beta}-2\omega_{\alpha})$, we see that the only weight having a non-trivial cohomology is equal to $\omega_{\beta}-2\omega_{\alpha}=-\alpha$ and $\mathop{\mathrm{R\Gamma}}\nolimits(X,\mathcal{L}(-\alpha))=\sf k[-1]$. Coming back to the sequence (96) and remembering the shift by $[3]$, we obtain the isomorphism $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\mathcal{E})=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf S}\otimes\mathcal{L}(\rho)[-3],\Lambda^{2}\mathcal{U}_{2}^{\perp})=\sf k$ and hence the required isomorphism in 6). ∎
@@ 14.16 Type ${\bf G}_{2}$ over ${\mathbb{Z}}$
The non-vanishing statements of Lemmas 14.31, 14.35, and 14.40 boiled down to non-vanishing cohomology groups of line bundles either in degree zero, or in the top degree, or in degree one. By [3, Corollary 2.6] and [29, II, Section 8.8, (6)], those cohomology groups are free $\mathbb{Z}$-modules considered as cohomology groups of line bundles over $\mathbb{G}/\mathbb{B}$. Thus, the results of the above lemmas hold over ${\mathbb{Z}}$.
14.16.1 Type ${\bf G}_{2}$ over ${\mathbb{Z}}$, without proofs
Let $\alpha$ be the short simple root, $\beta$ the long one. Put $a=s_{\alpha}$, $b=s_{\beta}$. The exceptional $\mathbb{B}$-modules $X_{v}$ have no ${\mathbb{Z}}$ torsion. Over ${\mathbb{Z}}[1/2]$ each $X_{v}$ has a relative Schubert filtration and over ${\mathbb{Z}}[1/6]$ each $X_{v}^{*}$ has an excellent filtration. If $v,w\in W$ are distinct and $\ell(v)\leq\ell(w)$, then $\mathop{\mathrm{RHom}}\nolimits^{\bullet}(\mathcal{L}(X_{v}),\mathcal{L}(X_{w}))=0$.
The natural map $\phi_{w}:P(-e_{w})^{*}\to X_{w}$ is injective for $w\in W$.
Some properties of ${\mathsf{coker}}(\phi_{w})$:
${\mathsf{coker}}(\phi_{\mathrm{id}})=0$.
${\mathsf{coker}}(\phi_{a})\cong{\mathbb{Z}}/2{\mathbb{Z}}$.
${\mathsf{coker}}(\phi_{b})\cong\nabla(\omega_{\alpha})/3\nabla(\omega_{\alpha})$.
$0\to\Delta(\omega_{\alpha})\to{\mathsf{coker}}(\phi_{ab})\to{\mathbb{Z}}\to 0$ is exact.
${\mathsf{coker}}(\phi_{ba})\cong{\mathbb{Z}}/2{\mathbb{Z}}$.
$0\to P(-e_{ba})^{*}\to{\mathsf{coker}}(\phi_{bab})\to{\mathbb{Z}}\to 0$ is exact.
${\mathsf{coker}}(\phi_{aba})\cong{\mathbb{Z}}$.
$0\to P(-e_{aba})^{*}\to{\mathsf{coker}}(\phi_{abab})\to{\mathbb{Z}}\to 0$ is exact.
${\mathsf{coker}}(\phi_{baba})={\mathsf{coker}}(\phi_{babab})={\mathsf{coker}}(\phi_{ababa})={\mathsf{coker}}(\phi_{bababa})=0$.

### v3

@@ 14.14 Type ${\bf G}_{2}$
in which the bundles from the block ${{\sf B}_{i}},i=0,\dots,6$ are shifted by $[-i]$ in the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}_{2}/{\mathbf{B}})$. In other words, the above collection is obtained from the collection Lemma 14.38 by tensoring with $\mathcal{L}(\rho)$ and applying the shift functor relative to the corresponding block. Then the collection above is right dual to that of Lemma 14.38.
@@ 14.16 Type ${\bf G}_{2}$ over ${\mathbb{Z}}$
The non-vanishing statements of Lemmas 14.32, 14.36, and 14.41 boiled down to non-vanishing cohomology groups of line bundles either in degree zero, or in the top degree, or in degree one. By [5, Corollary 2.6] and [32, II, Section 8.8, (6)], those cohomology groups are free $\mathbb{Z}$-modules considered as cohomology groups of line bundles over $\mathbb{G}/\mathbb{B}$. Thus, the results of the above lemmas hold over ${\mathbb{Z}}$.
15 The left dual exceptional collection to the $X_{v}$’s, conjecturally
Recall the notation ${\mathcal{P}}_{v}=\mathcal{L}(P(-e_{v}))$ and ${\mathcal{Q}}_{v}=\mathcal{L}(Q(e_{v}))$ for $v\in W$. Some evidence suggests that the following could be true:
$R\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{P}}_{v}\otimes{\mathcal{Q}}_{w_{0}v}\otimes\mathcal{L}(-\rho))={\sf k}[-\ell(w_{0}v)].$ (100)
We have in any case, provided $0\in\overline{C}_{\mathbb{Z}}$,
$R\Gamma(\mathbf{G}/\mathbf{B},\mathcal{L}(-e_{v})\otimes\mathcal{L}(e_{w_{0}v})\otimes\mathcal{L}(-\rho))={\sf k}[-\ell(w_{0}v)].$ (101)
This uses that $-e_{v}+e_{w_{0}v}-\rho=(w_{0}v)^{-1}\rho-\rho$ and Theorem 14.1.
There is also evidence that
$R\Gamma({\mathbf{G}/\mathbf{B}},{\mathcal{Q}}_{v}^{\ast}\otimes\mathcal{L}(-\rho)\otimes{\mathcal{Q}}_{w})\mbox{ vanishes unless }w_{0}v\geq w,$ (102)
and
$R\Gamma({\mathbf{G}/\mathbf{B}},{\mathcal{Q}}_{v}^{\ast}\otimes\mathcal{L}(-\rho)\otimes{\mathcal{Q}}_{w_{0}v}))={\sf k}[-\ell(w_{0}v)],$ (103)
and
$R\Gamma({\mathbf{G}/\mathbf{B}},{\mathcal{P}}_{v}\otimes\mathcal{L}(-\rho)\otimes{\mathcal{P}}_{w}^{\ast})\mbox{ vanishes unless }w_{0}v\leq w,$ (104)
and
$R\Gamma({\mathbf{G}/\mathbf{B}},{\mathcal{P}}_{v}\otimes\mathcal{L}(-\rho)\otimes{\mathcal{P}}_{w_{0}v}^{\ast})={\sf k}[-\ell(w_{0}v)].$ (105)
Assuming (102), (103), (104), (105), it is not difficult to see with equation (33) that $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}({\mathcal{X}}_{v}\otimes\mathcal{L}(\rho),{\mathcal{X}}_{w})$ vanishes unless $w_{0}v=w$. And that $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}({\mathcal{X}}_{w_{0}v}\otimes\mathcal{L}(-\rho),{\mathcal{X}}_{v}))={\sf k}[-\ell(w_{0}v)]$.
We can now state:
Conjecture 15.1.
The left dual exceptional collection to the ${\mathcal{X}}_{v}$’s consists of the ${\mathcal{X}}_{w_{0}v}\otimes\mathcal{L}(\rho)[-\ell(w_{0}v)]$.